OpenStudy (anonymous):

@FoolForMath Can you help me? Suppose that a and b vary inversely and that b=5/6 when a=3 Write a function that models the inverse variation and find b when a = 8

OpenStudy (anonymous):

Well, b=k/a. So 5/6=k/3. That makes k=(5)(3)/(6). Simplify this. This k will then be your proportionality constant. Therefore, we can look for b when a=8 since the only unknown would just be b and not k also. b=k/8. Just substitute and evaluate, and you have your answer.

OpenStudy (anonymous):

The basic form of inverse proportion is \[a=k/b\] First you need to find what the proportionality constant k is. Just insert the value of a and b given in the problem. \[3=k/(5/6) = 6k/5\] \[15=6k\] \[k=2.5\] Once you find k, then the relationship between a and b is known. Just substitute a with 8 to get your final answer :) \[8=k/b\] \[b=2.5/8\] \[b=5/16\]

OpenStudy (anonymous):

i got 5/6a;5/16 i'm right?

OpenStudy (anonymous):

or is it 5/2a;5/16...?

OpenStudy (anonymous):

a=5/2b or b=5/2a yes 5/16 is the correct answer

OpenStudy (anonymous):

okay, thank you guys SO much. I really appreciate you guys!